OpenStudy (anonymous):

solve -11≤3x+1 and 3x+1>13

OpenStudy (anonymous):

\[-11\leq 3x+1\] \[-12\leq 3x\] \[-4\leq x\] is the first one.

OpenStudy (anonymous):

\[3x+1\geq 13\] \[3x\geq 12\] \[x\geq 4\] is the second

OpenStudy (anonymous):

now if it really says solve -11≤3x+1 AND 3x+1>13 there is no solution since there is nothing that is both bigger then 4 and at the same time less than -4 if it says 11≤3x+1 OR 3x+1>13 then the solution is \[(-\infty,-4)\cup (4,\infty)\]